OpenStudy (phebe):

HELPP!!!! WANNNTTEDD!!! What is the simplified form of 3 root5c*root15n^9?

OpenStudy (phebe):

@ineedhelphaha134

OpenStudy (phebe):

@wolfe8

OpenStudy (phebe):

want answer choices

OpenStudy (wolfe8):

O.o I'm not sure which is the root which is the number here

OpenStudy (phebe):

3 its before the root

OpenStudy (wolfe8):

Ok so is it \[\sqrt[5]{3}c \sqrt[15]{n ^{9}}\] or what?

OpenStudy (phebe):

no

OpenStudy (phebe):

its 3squared5c*squared15c^3

OpenStudy (wolfe8):

That makes no sense to me. Is it \[3^{2}5c ^{2}15c ^{3}\] ?

OpenStudy (phebe):

im doing square root

OpenStudy (wolfe8):

I honestly don't understand how the question is

OpenStudy (phebe):

ok hold on k

OpenStudy (phebe):

\[\sqrt[3]5c{?}*\sqrt{15c^3?}\]

OpenStudy (phebe):

there's the question rit there @wolfe8

OpenStudy (wolfe8):

Ah ok so you can write roots as 1 over the root. So square root is to the power of 1/2, cube root is 1/3, and so on.

OpenStudy (phebe):

k so u think u can help me

OpenStudy (wolfe8):

Yea I just did. We can only give hints and help on OS. Try what I said.

OpenStudy (phebe):

ohh well im in a hurry so can you give me the answer

OpenStudy (anonymous):

pebe go here http://openstudy.com/study#/updates/4f832000e4b0505bf08429e4 already asked once this question has been

OpenStudy (wolfe8):

Nope. Not from me. Sorry that's against the rules

OpenStudy (phebe):

ok thanx @dajiawestlake

OpenStudy (anonymous):

no problem it shows the steps there and the other part.... also its DaJokerBOIIII

OpenStudy (anonymous):

do I get a medal lol?